Please find the Degree of unsaturation, IR analysis, H-NMR, and unknown identity. Thanks

Please find the Degree of unsaturation, IR analysi
Answer #1

Solution :-

C3H6O2

Hydrocarbon with the 3 carbon has 8 hydrogens

So the number of hydrogen deficiency = 8-6 = 2

Number of unsaturation = hydrogen defficieny /2

                                           = 2/2

                                           = 1

Now the IR analysis

In the IR spectrum it shows peak at the 1716 cm-1 this means the compound have carbonyl group (C=O)

Also it shows the broad peaks at about 3200